OpenStudy (anonymous):

complete the square: x^2+4x

OpenStudy (anonymous):

To complete the square you must first make sure the coefficient of x^2 is 1, and if it is, then all you have to do is take 1/2 of the coefficient of the x term, in this case 4, and after taking 1/2 of 4, you square that result. This result is then added to the expression creating your perfect square.

OpenStudy (anonymous):

\[x^2+ 4x+( \frac{4}{2})^2=x^2+ 4x+ 4\]

OpenStudy (anonymous):

Exactly as Luis has demonstrated.

OpenStudy (anonymous):

it has the answer set up like: (x+___)^2 + (___) ?

OpenStudy (anonymous):

ok, so factor the result that Luis has, \[x ^{2}+4x+4=(x+2)^{2}+(-4)\]

OpenStudy (anonymous):

You added 4 to the equation so you must subtract 4 from the erquation to keep it balanced.

OpenStudy (anonymous):

equation*

OpenStudy (anonymous):

thank you so much @calmat01 ! how would I graph that?

OpenStudy (anonymous):

Ok, so your equation now looks like this \[y=(x+2)^{2}-4\] This is a parabola whose vertex occurs at the values x=-2 y=-4.

OpenStudy (anonymous):

x=-2 comes from replacing x with -2 in the equation to make the expression in the parenthesis zero, and if what is inside is zero, then y must be -4.

OpenStudy (anonymous):

Now once you have the vertex, just replace the x with two more values, one to the left of x=-2 and one to the right of x=-2.;

OpenStudy (anonymous):

well it would be shifted left 2 units right and then down 4 units

OpenStudy (anonymous):

That is correct. It's vertex is (-2, -4) so, just pick one more value to the left of x=-2 and one more value to the right of x=-2.

OpenStudy (anonymous):

so would the parabola be facing up or down?

OpenStudy (anonymous):

Can you answer that question based on our equation?

OpenStudy (anonymous):

up because it's positive right

OpenStudy (anonymous):

Exactly, the coefficient of the squared term is positive, so the graph points up. Good!

OpenStudy (anonymous):

thank you so much for all of your help:)

OpenStudy (anonymous):

No Prob. Glad I could help!